# sample.bib

### v4

@@ 4 Experimental Evaluation
We designed experiments to accomplish two primary objectives: (a) to demonstrate the success of our RATE model in memory-intensive environments (T-Maze [ni2023when], ViZDoom-Two-Colors [memup], Memory Maze [pasukonis2022memmaze], Minigrid.Memory [MinigridMiniworld23]), and (b) to examine the versatility of the proposed model (Atari games [bellemare2013arcade] and MuJoCo control tasks [d4rl]). We chose Decision Transformer (DT) [chen2021decision] as the main baseline for comparison with RATE because it is an offline model that works in the sequence modeling paradigm for RL. It has had many branches [yamagata23dt, wu2024elastic, zheng22c], can handle both discrete and continuous action spaces, unlike TAP [jiang2023efficient], and has no memory mechanism.
@@ 4.1 Memory-intensive environments
    ViZDoom-Two-Colors [memup] – The agent in an acid-filled room observes a quickly disappearing green or red pillar. To stay alive, the agent must recall the pillar’s color and gather items of the same color.
@@
    T-Maze [ni2023when] – An agent navigates a T-shaped corridor, receiving a clue at the start about which direction to turn at the end of the corridor. The task tests memory in prolonged environments with sparse rewards, with the agent only receiving a reward at the end.
@@
    MemoryMaze [pasukonis2022memmaze] – An agent navigates a maze, seeking objects matching the color of its view frame. The frame color changes after each successful find. The goal is to collect the most matching objects within a time limit.
@@
    Minigrid.Memory [MinigridMiniworld23] – A similar task to T-Maze, but with different observation spaces and reward functions (see Appendix ATable 4). Another important difference is that the agent appears at a random point at the beginning of the episode, not at the beginning of the corridor. In the case of Minigrid.Memory it is necessary to reach that clue first (in T-Maze it is a memory problem, in Minigrid.Memory it is a memory and credit assignment problem [ni2023when]).
@@
For more information about environments, see Appendix A.
In each of the experiments, the RATE configurations were chosen such that the context length $K$ for the DT was equal to the effective context length $K_{eff}$ of the RATE. Thus, the actual context length for RATE was $N$ times smaller than the DT context length, where $N$ is the number of segments into which the trajectory is divided. At the same time, RATE and DT process the same trajectory fragments during training. More information about the training procedure for each environment can be found in Appendix C.
In all experiments, we use the same hyperparameters for RATE as for DT for each considered environment to simplify the comparison and listed them in Table 5. Note that the difference between the number of DT and RATE parameters is small: $\delta p$ = d_model $\times$ num_mem_tokens $\sim 10^{3}$.
Figure 3: Results for the ViZDoom-Two-Colors with (a) and without (b) pillar in the first 45 steps of the episode. (c) – results of inference as a function of the target return. The training dataset is balanced by the pillar’s color. We demonstrate mean $\pm$ standard error for 6 runs with 100 seeds.
@@ 4.2 Experimental Results
Figure 4: Inference results of RATE and DT at environment steps 0 – 89 (top) and 90 – 179 (bot).
@@
Figure 3 and  Figure 4 displays the results for RATE and DT. Results are presented for two configurations of inference: with the pillar visible at the beginning of the episode for the first $45$ steps and with the pillar turned off. As can be seen from the left part of the Figure 3, DT is not able to remember the color of the pillar, so the agent has a bias towards green objects (despite the fact that the dataset is balanced). This bias is due to the bias of the original A2C model used to generate the dataset – A2C model often switches to collecting green items when red items are to be collected. At the same time, RATE successfully remembers the color of the pillar, as reflected in the agent’s ability to collect not only green objects, but also red objects if required.
This conclusion of DT’s inability to use information out of context as opposed to RATE is supported by the experimental results presented in Figure 4. The top part of the Figure 4 shows the inference results for the first $90$ steps of the environment, when the information about pillar fits completely into the DT context and RATE effective context. In turn, the bottom part shows the inference results for the next $90$ environment steps, when the pillar color information starts to disappear from the DT context and RATE effective context (the context window moves as a sliding window). As a result, there is a drop in total reward for red-pillar environments by almost a factor of $2$, indicating DT’s inability to memorize information to use it out of context. In turn, for RATE, the values of total reward in the first and second cases are almost unchanged, indicating RATE’s ability to utilize information outside of the current context.
This fact indicates that RATE has indeed remembered the color of the pillar. Furthermore, even in just the first $90$ environment steps, divided into $N=3$ segments, RATE successfully learns the policy and has an average lifetime of $1700$ steps ($\sim 20$ times more than the number of steps used in training). It is worth noting that it is incorrect to compare RATE and A2C, since RATE is trained offline and A2C is trained online. Figure 3 (c) was constructed to investigate the dependence of model performance on the target reward fed at output. In this paper, we used an empirical estimate of the target reward as the average of the top-10 total rewards in the training dataset.
@@
The Figure 5 (a) shows the results of the DT and RATE inference. The inference of these models trained on trajectories of lengths $90$ and $150$ was conducted on corridor sizes 30303030900900900900, i.e., much larger than the context length $K=30$ of the models. The figure demonstrates that DT’s ability to solve the task is limited by the length of the training data’s context. Specifically, DT-5, trained on trajectories of length $5\times 30=150$, exhibits a significant drop in performance (reaching a random turn guessing accuracy of $0.5$) when tasked with inference on corridors exceeding $150$ in length. This proofs that DT-5 struggles to generalize beyond the context it was trained on.
Figure 5: Results for the T-Maze (a, b, c) and Minigrid.Memory (d).
In turn, RATE-3 and RATE-5 (trained on $3\times 30=90$ and $5\times 30=150$ steps, respectively) perform significantly better at inference corridor lengths longer than the model saw during training. This confirms the ability of the RATE model to successfully memorize important information and carry it through a long time.
Table 1: Raw scores for Atari games for DT and RATE. Results are presented in terms of the mean of 3 runs across 100 seeds $\pm$ standard deviation. Green – top-1 result, light green – top-2 result within the standard deviation.
Environment DT DMamba MambaDM RATE
Breakout 76.9 $\pm$ 27.3 70.6 $\pm$ 9.3 106.9 $\pm$ 5.8 119.8 $\pm$ 15.9
Qbert 2215.8 $\pm$ 1523.7 5786.0 $\pm$ 1295.2 10052.5 $\pm$ 1116.5 12931.4 $\pm$ 820.9
SeaQuest 1129.3 $\pm$ 189.0 992.1 $\pm$ 57.7 1286.0 $\pm$ 42.0 1013.3 $\pm$ 34.6
Pong 17.1 $\pm$ 2.9 1.6 $\pm$ 15.3 18.4 $\pm$ 0.8 18.7 $\pm$ 0.5
Minigrid.Memory.
The Figure 5 (d) shows the results of RATE and DT inference in the Minigrid.Memory in mazes of larger size than in the training data, with the context size unchanged. Unlike the previously discussed T-Maze, the credit assignment problem is also solved here, since the agent first has to reach the oracle and find out which object to turn towards in the future. As shown in the Figure 5 (d), RATE significantly outperforms DT for inference at any length.
@@
Despite the lack of policy in the training data, even under the assumption that offline learning does not train the agent to learn exploration, it was possible to obtain better metrics for RATE than for DT. Thus, in terms of mean value $\pm$ standard error of the mean, the results obtained for $1000$ steps of the environment for RATE are $1.91\pm 0.10$ and for DT are $1.45\pm 0.23$. Results are shown for $3$ model runs at inference at $100$ seeds.
The results for Atari and MuJoCo are presented in Table 1 and Table 2. Results for DMamba [ota2024decision] and MambaDM [cao2024mamba] are from the corresponding papers. A more detailed description of the results obtained can be found in the Appendix C. The authors of MambaDM [cao2024mamba] have not published code at this time, so experiments on memory-intensive environments were not performed.
@@
Table 2: Scores normalized according to the protocol in [d4rl] for MuJoCo control tasks. ME – Medium-Expert dataset, M – Medium dataset, MR – Medium-Replay dataset. RATE outperforms DT in 9/9 of the cases. Green – top-1 result, light green – top-2 result within the standard deviation.
@@
Dataset Environment CQL DT TAP DMamba MambaDM RATE
ME HalfCheetah 91.6 86.8 $\pm$ 1.3 91.8 $\pm$ 0.8 91.9 $\pm$ 0.6 86.5 $\pm$ 1.2 87.4 $\pm$ 0.1
ME Hopper 105.4 107.6 $\pm$ 1.8 105.5 $\pm$ 1.7 $1$$\pm$ 0.3 110.5 $\pm$ 0.3 112.5 $\pm$ 0.2
ME Walker2d 108.8 108.1 $\pm$ 0.2 107.4 $\pm$ 0.9 108.3 $\pm$ 0.5 108.8 $\pm$ 0.1 108.7 $\pm$ 0.5
M HalfCheetah 44.4 42.6 $\pm$ 0.1 45.0 $\pm$ 0.1 42.8 $\pm$ 0.1 42.8 $\pm$ 0.1 43.5 $\pm$ 0.3
M Hopper 58.0 67.6 $\pm$ 1.0 63.4 $\pm$ 1.4 83.5 $\pm$ 12.5 85.7 $\pm$ 7.8 77.4 $\pm$ 1.4
M Walker2d 72.5 74.0 $\pm$ 1.4 64.9 $\pm$ 2.1 78.2 $\pm$ 0.6 78.2 $\pm$ 0.6 80.7 $\pm$ 0.7
MR HalfCheetah 45.5 36.6 $\pm$ 0.8 40.8 $\pm$ 0.6 39.6 $\pm$ 0.1 39.1 $\pm$ 0.1 39.0 $\pm$ 0.6
MR Hopper 95.0 82.7 $\pm$ 7.0 87.3 $\pm$ 2.3 82.6 $\pm$ 4.6 86.1 $\pm$ 2.5 83.7 $\pm$ 8.2
MR Walker2d 77.2 66.6 $\pm$ 3.0 66.8 $\pm$ 3.1 70.9 $\pm$ 4.3 73.4 $\pm$ 2.6 73.7 $\pm$ 1.4
Action Associative Retrieval.
As shown above, in experiments with DT in the T-Maze environment, the model has a Success Rate of $50\%$ for inference at corridor
Figure 6: Action Associative Retrieval environment.
lengths longer than the transformer context length. This is due to the fact that even a DT trained on balanced data has a slight bias in the predicted probability towards one of the two required actions, which leads to the fact that when $t>K$ the agent constantly produces only one target junction action. In turn, the presence of memory in the agent allows us to combat this problem. To test whether an agent has memory in a more explicit way, we design an Action Associative Retrieval (AARFigure 6Appendix B) environment.
where $a^{0},a^{1}\in\mathcal{A}$ two mutually exclusive actions leading to a reward; $t$ is the step at which the final action is required; $N_{0},N_{1}$ are the number of experiments in environments where target action $a_{t}=a^{0}$ and $a_{t}=a^{1}$, respectively.
The results of experiments in the AAR environment are presented in Figure 7. The first $1\%$ of training steps was removed because it corresponds to the beginning of the training and is unrepresentative. Blue dots correspond to the beginning of training, red dots to the end of training. As can be seen from Figure 7, during training, the probabilities $p_{i}(a_{t}=a^{0}|a_{0}=a^{0})$ and $p_{i}(a_{t}=a^{1}|a_{0}=a^{1})$ on the training trajectories have a strong positive correlation ($R_{train}^{DT}=1.00$ and $R_{train}^{RATE}=0.85$), where $R$ correlation coefficient. This indicates that within-context (effective context) DT and RATE models are able to predict both $a^{0}$ and $a^{1}$ actions equally well.
At the same time, during validation, for the RATE model this pattern is preserved - the red points corresponding to the probabilities of choosing actions $a^{0}$ and $a^{1}$ are in the upper right part of the graph, positive correlation persists ($R_{val}^{RATE}=0.71$). On the other hand, in the DT case, the red cluster of points is shifted towards the probability of choosing action $a^{1}$ even when action $a^{0}$ is required to be chosen. Thus, in sum, these probabilities are less or equal to $1$, as evidenced by a strong negative correlation ($R_{val}^{DT}=-0.94$). The results confirm the inability of DT to generalize on trajectories whose lengths exceed the context length and the ability of RATE to handle such tasks.

### v5

@@ 4 Experimental Evaluation
We designed experiments to accomplish two primary objectives: (a) to demonstrate the advantage of our RATE model in memory-intensive environments (T-Maze (Ni et al., 2023), ViZDoom-Two-Colors (Sorokin et al., 2022), Memory Maze (Pasukonis et al., 2022), Minigrid-Memory (Chevalier-Boisvert et al., 2023)), and (b) to investigate the effectiveness of the proposed model in classical MDPs to demonstrate its versatility (Atari (Bellemare et al., 2013) and MuJoCo (Fu et al., 2021)).
For comparison with RATE, we chose DT (Chen et al., 2021) as the main baseline, and adapted the memory-augmented architectures Recurrent Memory Transformer (RMT) (Bulatov et al., 2022) and Transformer-XL (TrXL) (Dai et al., 2019) developed for NLP tasks to the RL domain. Information about the environments used can be found in Table 7.
@@ 4.1 Memory-intensive environments
To test RATE memory mechanisms, we use memory-intensive environments Figure 2, i.e., environments where the agent requires memory to operate successfully. A brief description of these environments is presented below, and a full description and data collection methodology can be found in the Appendix B.
@@
    ViZDoom-Two-Colors (Sorokin et al., 2022) – an agent in an acid-filled room observes a quickly disappearing green or red pillar. To stay alive, the agent must recall the pillar’s color and gather items of the same color.
@@
    T-Maze (Ni et al., 2023) – an agent navigates a T-shaped corridor, receiving a clue at the start about which direction to turn at the end of the corridor. The task tests memory in a prolonged sparse reward environment (the agent receives a reward only at the end).
@@
    Memory Maze (Pasukonis et al., 2022) – an agent navigates a maze, seeking objects matching the color of its view frame. The frame color changes after each successful find. The goal is to collect the most matching objects within a time limit.
@@
    Minigrid-Memory (Chevalier-Boisvert et al., 2023) – a similar task to T-Maze, but with different observation spaces and reward functions (see Appendix BTable 7). Another important difference is that the agent appears at a random point at the beginning of the episode, not at the beginning of the corridor. Thus, in the case of Minigrid-Memory, it is necessary to reach that clue first (in T-Maze it is a memory problem, in Minigrid-Memory it is a memory and credit assignment problem (Ni et al., 2023)).
@@
In the experiments, the same hyperparameters presented in Table 8 were used for RATE, DT, RMT, and TrXL to simplify the comparison. The context length $K$ and the number of segments $N$ were chosen so that the effective context $K_{eff}=K\times N$ for RATE, RMT, and TrXL covers important events in memory-intensive environments during training. In turn, since DT has no memory mechanisms, for it $K=K_{eff}$. Thus, the context length $K$ for RATE, RMT and TrXL is less than the context length for DT by a factor of $N$, but all models process the same parts of trajectories. More information about the training procedure for each environment can be found in Appendix D.
@@ 4.2 Experimental Results
In this section, the main experimental results for each of the environments are presented in the corresponding paragraphs. Additional results can be found in the Appendix F. For each of the experiments, the same techniques were used for all models to obtain the results presented in the Appendix E. Unless otherwise indicated, all baselines were trained from scratch.
@@
The dataset for this environment was collected using a pre-trained Advantage Actor Critic (A2C) (Beeching et al., 2019), which has a slight bias in favor of selecting green items even if red items are required. However, the dataset is balanced by the pillars colors. Figure 3 (a) shows the inference results on all pillars, separately only on red pillars and separately only on green pillars. As illustrated in the Figure 3 (a), for inference with the presence of a disappearing pillar at the beginning of the episode, all baselines have an average total reward for inference on green pillars greater than for inference on red pillars.
To prove that this is not due to the peculiarities of the algorithms, but to the presence of bias in the data, we performed an additional inference without a pillar at the beginning of the episode, demonstrating the ability of all baselines to collect exclusively green items. As can be seen from Figure 3 (b), DT learns the distribution of training data and is unable to remember the pillar color. In turn, baselines with memory mechanisms such as RATE, RMT and TrXL are successful in this task. The poor performance of baselines with memory on red pillars without the pillar at the episode beginning proves that it is the color of the pillar that they remember.
Figure 3: Results for the ViZDoom-Two-Colors: with (a) and without (b) pillar in the first $45$ steps of the episode; calculated at environment steps 0089898989 (c) and 90909090179179179179 (d); depending on the return-to-go (e, f, g). Pillar disappears after first $45$ steps, $K_{eff}=90$.
This conclusion of DT’s inability to use information out of context as opposed to RATE is supported by the experimental results presented in Figure 3 (c, d), which illustrates the inference results for the first $90$ steps, where the pillar are entirely captured in the effective context, and the subsequent $90$ steps, where the pillar color information begins to disappear from the effective context (the context window moves as a sliding window). As a result, there is a drop in total reward for red-pillar environments by almost a factor of two, indicating DT’s inability to memorize information to use it out of context. In turn, for RATE, RMT, and TrXL, the values of total reward in the first and second cases are almost unchanged, indicating their ability to utilize information outside of the current context window.
Figure 3 (e, f, g) demonstrates the dependence of model performance on the return-to-go. In this paper, we used an empirical estimate of the target reward as the average of the top-$10$ total rewards in the training dataset. As can be seen, RATE not only significantly outperforms DT, but also outperforms the memory-augmented models RMT and TrXL.
Furthermore, Table 1 demonstrates that RATE outperforms not only transformer-based models (RMT, TrXL), but also recurrent baselines, which forget the pillars color fairly quickly and start to collecting green items like DT.
@@
The Figure 4 shows the inference results for the T-Maze. To validate an agent’s long-term memory capabilities, we performed training on trajectories of length $90$ and inference on corridors of size 30303030900900900900, i.e. much larger than the effective context $K_{eff}=90$ of models. The Figure 4 demonstrates that DT’s ability to solve the task is limited by the length of the corridors in the training data. Specifically, DT-3, trained on trajectories of length $3\times 30=90$ with $K=K_{eff}=90$, exhibits a significant drop in performance (reaching a random turn guessing accuracy of $0.5$) when tasked with inference on corridors exceeding $90$ in length.
Figure 4: Results for the T-Maze environment with $K_{eff}=30\times 3=90$. The notations are represented as MODEL-N, where N is the number of segments into which the trajectories are divided.
In turn, RATE-3, RMT-3, and TrXL-3 (trained on $3\times 30=90$ steps) perform significantly better at inference corridor lengths longer than the model saw during training. Moreover, RATE-3 outperforms other memory-augmented baselines, indicating its ability to perform effectively in sparse environments. This confirms the ability of the RATE model to successfully memorize important information and retain it through a long time.
Additionally, Table 1 compares RATE with recurrent baselines. The results indicate that these baselines, unlike RATE, cannot handle sparse information, as shown by SR $=0.5$.
Table 1: Comparison of Transformer (DT), RNN (Decision LSTM (DLSTM) (Siebenborn et al., 2022), Decision GRU (DGRU)) and SSM (Decision Mamba (DMamba) (Gu & Dao, 2023)) models with RATE in memory-intensive environments. Results indicate the inability of RNN and SSM models to perform successfully in long-term memory sparse (T-Maze) and continuous (ViZDoom-Two-Colors) environments, in contrast to RATE. SR – Success Rate. DGRU is obtained by replacing the LSTM block with the GRU block (Chung et al., 2014) in DLSTM.
T-Maze DLSTM (Siebenborn et al., 2022) ($K=90$) DGRU ($K=90$) DMamba (Gu & Dao, 2023) ($K=90$) DT (Chen et al., 2021) ($K=90$) RATE ($K_{eff}=3\times 30=90$) SR ($T=90$) 0.5 $\pm$ 0.0 0.5 $\pm$ 0.0 0.5 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 ViZDoom-Two-Colors Reward[Total] 13.1 $\pm$ 0.6 12.9 $\pm$ 0.2 26.9 $\pm$ 1.9 24.8 $\pm$ 1.4 41.5 $\pm$ 1.0 Reward[Reds] 8.8 $\pm$ 0.7 9.4 $\pm$ 0.5 6.9 $\pm$ 0.4 7.2 $\pm$ 0.4 38.2 $\pm$ 5.1 Reward[Greens] 17.5 $\pm$ 1.6 16.3 $\pm$ 0.8 46.9 $\pm$ 4.2 42.3 $\pm$ 3.3 44.7 $\pm$ 5.8
Figure 5: Results for the Minigrid-Memory environment, $K_{eff}=10\times 3=30$.
Minigrid-Memory.
The Figure 5 shows the results for the Minigrid-Memory. Training was conducted on grids of size $31$x31313131, inference was conducted on grids of size $11$x1111111191919191x91919191. Unlike the previously discussed T-Maze, the credit assignment problem is also addressed here, since the agent first has to reach the oracle and find out which object to turn towards in the future.
The results show that TrXL performs better than the other baselines on grids of size smaller or equal to those used in training, and worse than the other baselines on larger grids, that is, it interpolates well and extrapolates poorly. In turn, for RMT we observe exactly the opposite situation: RMT interpolates poorly and extrapolates well. RATE performs slightly worse than TrXL but better than RMT on small grid sizes, and slightly worse than RMT but better than TrXL on large grid sizes, but on average has interpolation and extrapolation abilities better than RMT and TrXL individually, as well as greater stability.
@@
Table 2 shows the results of comparing RATE with other basic baselines with and without memory. The results show that RATE is able to memorize the implicit information like maze structure more efficiently, which is reflected in a higher average reward per episode.
Table 2: Results for the Memory Maze $9$x9999 environment.
DT (Chen et al., 2021) RMT (Bulatov et al., 2022) TrXL (Dai et al., 2019) RATE Reward 6.83 $\pm$ 0.51 7.27 $\pm$ 0.21 7.12 $\pm$ 0.24 7.64 $\pm$ 0.41
Atari and MuJoCo.
The results for Atari and MuJoCo are presented in Table 3 and Table 4. Results for Decision Mamba (DMamba) (Ota, 2024) and Mamba as Decision Maker (MambaDM) (Cao et al., 2024) are from the corresponding papers. A more detailed description of the results obtained can be found in the Appendix D. The results demonstrate that RATE not only performs as well as the algorithms specifically designed for offline RL, but in many cases outperforms them in classical environments that do not require memory, which indicates the versatility of the model.
Table 3: Raw scores for Atari games. Green – top-1 result, light green – top-2 result within the standard deviation.
Environment CQL (Kumar et al., 2020) DT (Chen et al., 2021) DMamba (Ota, 2024) MambaDM (Cao et al., 2024) RATE Breakout 62.5 76.9 $\pm$ 27.3 70.6 $\pm$ 9.3 106.9 $\pm$ 5.8 111.0 $\pm$ 2.9 Qbert 14013.2 2215.8 $\pm$ 1523.7   5786.0 $\pm$ 1295.2     10052.5 $\pm$ 1116.5 12486.9 $\pm$ 280.4 SeaQuest 782.2 1129.3 $\pm$ 189.0 992.1 $\pm$ 57.7 1286.0 $\pm$ 42.0 1037.9 $\pm$ 53.7 Pong 18.8 17.1 $\pm$ 2.9     1.6 $\pm$ 15.3    18.4 $\pm$ 0.8    18.8 $\pm$ 0.3
@@
Table 4: Scores normalized according to the protocol in Fu et al. (2021) for MuJoCo control tasks. ME – Medium-Expert dataset, M – Medium dataset, MR – Medium-Replay dataset. RATE outperforms DT in 9/9 of the cases. Green – top-1 result, light green – top-2 result within the standard deviation.
@@
Dataset Environment CQL (Kumar et al., 2020) DT (Chen et al., 2021) TAP (Jiang et al., 2023) DMamba (Ota, 2024) MambaDM (Cao et al., 2024) RATE ME HalfCheetah 91.6 86.8 $\pm$ 1.3 91.8 $\pm$ 0.8 91.9 $\pm$ 0.6 86.5 $\pm$ 1.2 87.4 $\pm$ 0.1 ME Hopper 105.4 107.6 $\pm$ 1.8 105.5 $\pm$ 1.7 $1$$\pm$ 0.3 110.5 $\pm$ 0.3 112.5 $\pm$ 0.2 ME Walker2d 108.8 108.1 $\pm$ 0.2 107.4 $\pm$ 0.9 108.3 $\pm$ 0.5 108.8 $\pm$ 0.1 108.7 $\pm$ 0.5 M HalfCheetah 44.4  42.6 $\pm$ 0.1 45.0 $\pm$ 0.1   42.8 $\pm$ 0.1   42.8 $\pm$ 0.1 43.5 $\pm$ 0.3 M Hopper 58.0  67.6 $\pm$ 1.0   63.4 $\pm$ 1.4    83.5 $\pm$ 12.5 85.7 $\pm$ 7.8 77.4 $\pm$ 1.4 M Walker2d 72.5  74.0 $\pm$ 1.4   64.9 $\pm$ 2.1   78.2 $\pm$ 0.6   78.2 $\pm$ 0.6 80.7 $\pm$ 0.7 MR HalfCheetah 45.5  36.6 $\pm$ 0.8   40.8 $\pm$ 0.6   39.6 $\pm$ 0.1   39.1 $\pm$ 0.1   39.0 $\pm$ 0.6 MR Hopper 95.0 82.7 $\pm$ 7.0   87.3 $\pm$ 2.3 82.6 $\pm$ 4.6 86.1 $\pm$ 2.5 83.7 $\pm$ 8.2 MR Walker2d 77.2  66.6 $\pm$ 3.0   66.8 $\pm$ 3.1   70.9 $\pm$ 4.3 73.4 $\pm$ 2.6 73.7 $\pm$ 1.4 Mean 77.6 74.7 74.8 78.8 79.0 78.5